# The role of charm and unflavored mesons in prompt atmospheric lepton fluxes

### v1

@@ Abstract
The all-sky very-high-energy ($10^{4}-10^{6}$ GeV) atmospheric muon flux measured by IceCube shows a spectral hardening at the highest energies, indicating the presence of a prompt component. IceCube has also measured the atmospheric muon neutrino flux at high energy. However, since this flux is dominated by astrophysical neutrinos, only an upper bound can be placed on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ contribution. In this work, we provide a new evaluation of the prompt atmospheric muon flux including an intrinsic charm component in the cosmic ray-air interactions. The latter enhances the forward production of $\bar{D}^{0}$, $D^{-}$, and $\Lambda_{c}$, which subsequently decay into final states containing muons and muon neutrinos. We show how the increase in the prompt muon flux due to intrinsic charm is accompanied by a corresponding enhancement in the prompt muon neutrino flux. We implement different intrinsic charm production models in MCEq to calculate the resulting lepton fluxes. We discuss the challenges of achieving predictions that are simultaneously consistent with both IceCube’s high-energy atmospheric muon flux measurements and IceCube upper bound on the prompt muon neutrino flux, and we quantify the resulting discrepancies. As possible solutions, we explore scaling of the unflavored meson contributions to the prompt atmospheric muon flux to assess how such adjustments can reconcile these differences. The tensions emphasized in our work call for a refinement of the hadronic interaction models, especially the production of unflavored mesons, and for new experimental data sensitive to unflavored meson and heavy flavor production with reliable estimates of the associated uncertainties. We suggest that the energy and zenith angle dependence of muon and neutrino flux ratios from future neutrino telescope measurements may help to disentangle different scenarios.
@@ 1 Introduction
When high-energy cosmic rays enter the Earth’s atmosphere, they interact with air nuclei and initiate a cascade of secondary particles, including protons, neutrons, pions, kaons, and charm hadrons. These secondary hadrons undergo further interactions and decays, giving rise to the atmospheric fluxes of muons and neutrinos (see, e.g., ref. Gaisser:2002jj and references therein). Since the primary cosmic ray spectrum spans many orders of magnitude in energy, the secondary particles are also produced in a wide energy range Gaisser:2016uoy ; Lipari:1993hd . Among these, high-energy atmospheric leptons ($E>10^{4}$ GeV), in particular muons and neutrinos, are of central importance in astroparticle physics. The latter form a major background for neutrino telescopes searching for astrophysical neutrinos IceCube:2013low ; IceCube:2016umi ; IceCube:2020acn ; IceCube:2024fxo ; IceCube:2025dlr , making precise knowledge of the atmospheric neutrino fluxes essential. Such measurements, together with the complementary input from atmospheric muon fluxes, not only help testing and constraining hadronic interaction models (see, e.g., Fedynitch:2018cbl ; Fedynitch:2022vty ; Yanez:2023lsy ), but also enable robust searches for astrophysical neutrino signals.
@@
Atmospheric leptons are produced through a chain of interactions and decays: cosmic rays interact with air nuclei with average atomic mass number of $\langle A_{\rm air}\rangle=14.5$ to produce hadrons. Muons and neutrinos come from leptonic or semi-leptonic decays of these hadrons. To evaluate the atmospheric lepton flux, incident cosmic ray nuclei of mass number $A_{N}$, charge $Z$ and energy $E_{0}$ are treated in the superposition approximation, i.e., the nucleus is treated as an ensemble of $Z$ protons and $A_{N}-Z$ neutrons, each carrying energy $E_{0}/A_{N}$ Gaisser:2016uoy . The resulting all-nucleon cosmic ray spectrum at the top of the atmosphere is denoted by $\phi^{0}_{N}(E)$. Although the cosmic ray composition above $E_{0}\simeq 10^{5}-10^{6}$ GeV Albrecht:2021cxw has been measured only indirectly, there is reasonably good agreement among cosmic ray composition models that $\phi^{0}_{N}(E)\sim E^{-\gamma}$ with the value of $\gamma$ approximately constant but different in different energy ranges (e.g., broken power law) Gaisser:2011klf ; Gaisser:2013bla ; Hoerandel:2002yg .
@@
The magnitudes of the atmospheric lepton fluxes from hadron decays depend on the competition between hadron-air interactions and hadron decays. If the hadron interaction length is much longer than its decay length, the lepton flux at energy $E_{\ell}$ approximately follows the primary cosmic ray nucleon spectrum, scaling as $\sim E_{\ell}^{-\gamma}$. Conversely, when hadrons are more likely to interact than decay, the corresponding atmospheric lepton flux is suppressed, scaling as $E_{\ell}^{-\gamma-1}$ Lipari:1993hd . For the energies considered here, $E_{\ell}\gtrsim 10^{4}$ GeV, the contributions from $\pi^{\pm}$ and $K$ decays to muon and neutrino fluxes scale like $\sim E_{\ell}^{-\gamma-1}$. The combination of these two contributions is referred to as the conventional atmospheric lepton flux. In contrast, hadrons with much shorter lifetimes, such as charm hadrons, decay promptly over a wide energy range. The prompt atmospheric lepton flux from charm hadron decays scales as $\sim E_{\ell}^{-\gamma}$, up to lepton energies $E_{\ell}\sim 2\times 10^{7}$ GeV Gondolo:1995fq . In addition to charm hadron decays, prompt muon production also receives contributions from electromagnetic decays of light unflavored mesons such as $\eta,\ \eta^{\prime},\ \rho^{0},\ \omega,$ and $\phi$, despite their small branching fractions ($\sim 10^{-5}$$10^{-4}$) Illana:2010gh . Their extremely short lifetimes ensure decay before interaction, enhancing the prompt atmospheric muon flux at high energies by a factor of $\sim 2$ relative to the prompt neutrino flux.
@@
Figure 1: The atmospheric ($\mu^{+}+\mu^{-}$) flux and atmospheric ($\nu_{\mu}+\bar{\nu}_{\mu}$) flux, both scaled by $E^{3}$ and angle-averaged over zenith angles less than 60, from MCEq Fedynitch:2018cbl using H3a cosmic ray flux model Gaisser:2011klf and Sibyll-2.3c Riehn:2017mfm ; Fedynitch:2018cbl for interactions. The dashed lines show the prompt contributions, and the solid lines show the total contributions (conventional + prompt).
Prompt neutrinos from charm decays in the atmosphere are often considered an irreducible background to astrophysical neutrinos since both fluxes are nearly isotropic at Earth. Their spectral behaviors differ, offering possible means of discrimination between the prompt atmospheric background and astrophysical neutrinos. In this paper, we focus on using the atmospheric muon flux as a complementary probe to atmospheric neutrinos. While many previous studies have focused primarily on neutrino fluxes Laha:2016dri ; Halzen:2016thi ; Ostapchenko:2022thy ; Garzelli:2023jlq , our analysis expands to include both muons and neutrinos (see also, ref. Das:2025snq ). As we describe in the next section, measurements of the atmospheric muon flux, particularly in the energy range where the prompt muon flux begins to dominate, exceed theory predictions, although with large experimental error bars Soldin:2023lbr ; Soldin:2018vak ; IceCube:2015wro . To solve this discrepancy, we first consider potential contributions from intrinsic charm. The intrinsic charm contributions are also constrained by prompt neutrino measurements. Therefore, with existing upper bounds on the prompt atmospheric muon neutrino flux from IceCube IceCube:2016umi ; IceCube:2015wro , we also assess the extent to which the contribution of unflavored mesons to the prompt atmospheric muon flux may be enhanced.
The paper is organized as follows: in Section 2, we discuss the cascade equations that describe the development of the atmospheric lepton fluxes as cosmic rays and the hadrons they produce propagate through the atmosphere. We also show results from the cascade equations for muons and muon neutrinos, along with IceCube data and upper bounds on the prompt neutrino flux, respectively. In Section 3, we introduce the framework we use for incorporating intrinsic charm effects. Section 4 presents our results, followed by a discussion and our conclusions in Section 5. For completeness, the Appendix includes details of the charm particle decays to leptons, needed to evaluate the lepton fluxes.
@@ 2.1 Cascade equations
Quantitative evaluations of the atmospheric flux begin with the incident all-nucleon cosmic ray spectrum at the top of the atmosphere, $\phi_{N}^{0}(E)$. The flux per solid angle $\phi_{N}^{0}$ is isotropic at high energies. The cascade equations that describe the development of particle fluxes with subsequent interactions and decays Lipari:1993hd ; Gondolo:1995fq . The cascade equations are evaluated as a function of column depth $X$ which depends on the atmospheric density ($\rho$), path length ($l$) traversed by the particle in the atmosphere, altitude ($h$) and zenith angle ($\theta$),
@@
The cascade equations can be solved approximately in the hadron-decay-dominated and hadron-interaction-dominated regimes Lipari:1993hd ; Gondolo:1995fq . This method, called the $Z$-moment method, approximates eq. (3) for production and decay processes by putting $\phi_{k}(E^{\prime},X)/\phi_{k}(E,X)\simeq\phi_{k}(E^{\prime},0)/\phi_{k}(E,0)$ so that
@@
An alternative approach for evaluating the atmospheric lepton fluxes is to solve the cascade equations directly. The Matrix Cascade Equations (MCEq) code fedynitch2015calculationconventionalpromptlepton ; Fedynitch:2018cbl solves the cascade equations using a matrix approach. While the MCEq results are more precise, given hadronic production inputs, the $Z$-moment method is straightforward to implement. The two approaches differ by $5-20\%$ depending on energy and angle Gaisser:2019xlw . In this paper, we use the MCEq code with some modifications outlined below to evaluate all atmospheric lepton fluxes. Another approach is to use Monte Carlo air shower simulations. Recent work LudwigNeste:2025mcf shows good agreement with MCEq.
@@ 2.2 Muon and muon neutrino fluxes from MCEq
MCEq is equipped with several hadronic interaction models and cosmic ray nucleon flux models. We use the Sibyll2.3c Riehn:2017mfm ; Fedynitch:2018cbl model for hadronic interactions, and the H3a cosmic ray flux and composition model Gaisser:2011klf . We discuss the impact of different hadronic models and all-nucleon cosmic ray flux inputs in Section 4.3.
@@
The implementation of charm production in Sibyll-2.3c follows a phenomenological model based on soft and hard minijets Fedynitch:2018cbl . For LHCb kinematic distributions for charmed-meson production, the overall combination of perturbative and non-perturbative contributions in Sibyll-2.3c is consistent with the next-to-leading order QCD predictions of refs. Gauld:2015yia ; Garzelli:2016xmx ; PROSA:2015yid that come from collinear factorization with partonic cross sections from perturbative QCD and non-perturbative contributions from both parton distribution functions and fragmentation functions/hadronization. While some components of the implementation of charm production in Sibyll-2.3c are described as “intrinsic” charm, the charm hadrons do not have the same energy distribution, shape and normalization, as in the model for intrinsic charm that we introduce in Section 3. Even though Sibyll-2.3c’s implementation of charm production has a weighted mix of perturbative and non-perturbative components, we will nevertheless refer to Sibyll-2.3c’s contribution to the prompt atmospheric lepton fluxes as coming from perturbative QCD (pQCD($c\bar{c}$)) to distinguish it from the intrinsic charm contribution we introduce in Section 3.
@@
Figure 2 shows the MCEq results for the separate contributions of pions, kaons and (prompt) charm hadrons to the muon (left) and muon neutrino (right) atmospheric fluxes for zenith angle $\theta_{\rm zen}=0^{\circ}$ (vertical). As noted above, the muon flux also includes contributions from light unflavored mesons to its prompt component. The standard release of MCEq does not include $\Lambda_{c}$ production and decay. We include perturbative $\Lambda_{c}$ production by extracting the $D^{0}$ production spectrum from MCEq and applying a constant suppression factor of 0.10 that reflects the relative charm quark fragmentation fractions to $\Lambda_{c}$ compared to $D^{0}$ in proton-proton collisions Lisovyi:2015uqa . The final results of our paper are not very sensitive to this suppression factor because charm contributions to the atmospheric $\mu^{+}+\mu^{-}$, $\nu_{\mu}+\bar{\nu}_{\mu}$ and $\nu_{e}+\bar{\nu}_{e}$ fluxes are dominated by the semi-leptonic decays of $D^{\pm},\ D^{0}$ and $\bar{D}^{0}$ mesons. For semi-leptonic $\Lambda_{c}$ decays, we follow refs. Bugaev:1998bi and Buras:1976dg for the energy distributions of muons and neutrinos, which we have also described in the Appendix. While the details of the semi-leptonic $\Lambda_{c}$ decay distribution are not essential for the pQCD($c\bar{c})$ contributions to the prompt lepton fluxes, a $\Lambda_{c}$ is always produced in our implementation in intrinsic-charm induced events, as discussed in section 3.
@@
Figure 2 exhibits an high-energy behavior that reflects the dominance of the prompt contributions to the atmospheric lepton fluxes, visible in their harder spectra compared to the conventional fluxes. The plots show fluxes for vertical trajectories entering from the zenith, $\theta_{\rm zen}=0^{\circ}$. The zenith angle dependence is primarily in the conventional flux, where pions and kaons are more likely to interact than decay. The conventional flux increases with zenith angle according to $1/\cos\theta_{\rm zen}$ for $\theta_{\rm zen}\leq 70^{\circ}$ Klimushin:2000cy . At very high energies, the prompt contributions will also vary with zenith angle, however, at the energies we consider for the rest of the paper, $E\lesssim 10^{6}$ GeV, the prompt atmospheric lepton fluxes are essentially isotropic.
@@
We can also obtain di-muon pairs through the process $\gamma\to\mu^{+}\mu^{-}$. There are two possible sources for this channel. The first is from diffuse cosmic rays interacting in the atmosphere and producing high energy photons, which can subsequently convert into muon pairs. However, this contribution is expected to be very small even at the highest $\gamma$-ray energies Illana:2010gh . The second source is the diffuse $\gamma$-ray flux from the Galactic plane. According to ref. LHAASO:2023gne , at around 100 TeV the diffuse Galactic $\gamma$-ray flux is approximately $10^{-5}$ of the diffuse cosmic ray flux. Thus, diffuse Galactic $\gamma$-ray flux will also provide negligible contribution to the overall atmospheric muon flux.
Figure 3: Left: MCEq predictions for angle-averaged (up to zenith angle of $60^{\circ}$) conventional and prompt ($\mu^{+}+\mu^{-}$) flux, compared to the experimentally measured spectrum of high-energy muons using data for $\theta_{\rm zen}\leq 60^{\circ}$ from refs. Soldin:2023lbr ; Soldin:2018vak ; IceCube:2015wro . The unflavored and pQCD($c\bar{c}$) curves are nearly equal in the muon energy range shown in the figure. Right: Comparison of MCEq pQCD($c\bar{c}$) prompt ($\nu_{\mu}+\bar{\nu}_{\mu}$) flux with the upper limit on the prompt flux from the analysis of up-going muon tracks performed in Ref. IceCube:2016umi , which assumed a prompt neutrino energy spectrum with the same shape as in Enberg et al. PhysRevD.78.043005 . The angle-averaged (up to zenith angle of $60^{\circ}$) conventional neutrino flux from MCEq is also shown. The prompt neutrino flux is isotropic in the energy range considered here. The cosmic ray all-nucleon flux model H3a and the hadronic interaction model Sibyll-2.3c are used as input of MCEq.

### v2

@@ Abstract
The all-sky very-high-energy ($10^{4}-10^{6}$ GeV) atmospheric muon flux measured by IceCube shows a spectral hardening at the highest energies, indicating the presence of a prompt component. IceCube has also measured the atmospheric muon neutrino flux at high energy. However, since this flux is dominated by astrophysical neutrinos, only an upper bound can be placed on the prompt atmospheric $\nu_{\mu}+\bar{\nu}_{\mu}$ contribution. In this work, we provide a new evaluation of the prompt atmospheric muon flux including an intrinsic charm component in the cosmic ray-air interactions. The latter enhances the forward production of $\bar{D}^{0}$, $D^{-}$, and $\Lambda_{c}$, which subsequently decay into final states containing muons and muon neutrinos. We show how the increase in the prompt muon flux due to intrinsic charm is accompanied by a corresponding enhancement in the prompt muon neutrino flux. We implement different intrinsic charm production models in MCEq to calculate the resulting lepton fluxes. We discuss the challenges of achieving predictions that are simultaneously consistent with both IceCube’s high-energy atmospheric muon flux measurements and IceCube upper bounds on the prompt muon neutrino flux, and we quantify the resulting discrepancies. As possible solutions, we explore scaling of the unflavored meson contributions to the prompt atmospheric muon flux to assess how such adjustments can reconcile these differences. The tensions emphasized in our work call for a refinement of the hadronic interaction models, especially the production of unflavored mesons, and for new experimental data sensitive to unflavored meson and heavy flavor production with reliable estimates of the associated uncertainties. We suggest that the energy and zenith angle dependence of muon and neutrino flux ratios from future neutrino telescope measurements may help to disentangle different scenarios.
@@ 1 Introduction
When high-energy cosmic rays enter the Earth’s atmosphere, they interact with air nuclei and initiate a cascade of secondary particles, including protons, neutrons, pions, kaons, and charm hadrons. These secondary hadrons undergo further interactions and decays, giving rise to the atmospheric fluxes of muons and neutrinos (see, e.g., ref. Gaisser and Honda (2002) and references therein). Since the primary cosmic ray spectrum spans many orders of magnitude in energy, the secondary particles are also produced in a wide energy range Gaisser et al. (2016); Lipari (1993). Among these, high-energy atmospheric leptons ($E>10^{4}$ GeV), in particular muons and neutrinos, are of central importance in astroparticle physics. The latter form a major background for neutrino telescopes searching for astrophysical neutrinos Aartsen and others (2013a, 2016b, 2020); Abbasi and others (2024); Balagopal V. and others (2025); Abbasi and others (2025a), making precise knowledge of the atmospheric neutrino fluxes essential. Such measurements, together with the complementary input from atmospheric muon fluxes, not only help testing and constraining hadronic interaction models (see, e.g., Fedynitch et al. (2019); Fedynitch and Huber (2022); Yañez and Fedynitch (2023)), but also enable robust searches for astrophysical neutrino signals.
@@
Atmospheric leptons are produced through a chain of interactions and decays: cosmic rays interact with air nuclei with average atomic mass number of $\langle A_{\rm air}\rangle=14.5$ to produce hadrons. Muons and neutrinos come from leptonic or semi-leptonic decays of these hadrons. To evaluate the atmospheric lepton flux, incident cosmic ray nuclei of mass number $A_{N}$, charge $Z$ and energy $E_{0}$ are treated in the superposition approximation, i.e., the nucleus is treated as an ensemble of $Z$ protons and $A_{N}-Z$ neutrons, each carrying energy $E_{0}/A_{N}$ Gaisser et al. (2016). The resulting all-nucleon cosmic ray spectrum at the top of the atmosphere is denoted by $\phi^{0}_{N}(E)$. Although the cosmic ray composition above $E_{0}\simeq 10^{5}-10^{6}$ GeV has been measured only indirectly Kampert and Unger (2012), there is reasonably good agreement among cosmic ray composition models that $\phi^{0}_{N}(E)\sim E^{-\gamma}$ with the value of $\gamma$ approximately constant but different in different energy ranges (e.g., broken power law) Apel et al. (2012); Aab and others (2014); Rawlins (2020); Gaisser (2012); Gaisser et al. (2013); Hoerandel (2003); Dembinski et al. (2018).
@@
The magnitudes of the atmospheric lepton fluxes from hadron decays depend on the competition between hadron-air interactions and hadron decays. If the hadron interaction length is much longer than its decay length, the lepton flux at energy $E_{\ell}$ approximately follows the primary cosmic ray nucleon spectrum, scaling as $\sim E_{\ell}^{-\gamma}$. Conversely, when hadrons are more likely to interact than decay, the corresponding atmospheric lepton flux is suppressed, scaling as $E_{\ell}^{-\gamma-1}$ Lipari (1993). For the energies considered here, $E_{\ell}\gtrsim 10^{4}$ GeV, the contributions from $\pi^{\pm}$ and $K$ decays to muon and neutrino fluxes scale like $\sim E_{\ell}^{-\gamma-1}$. The combination of these two contributions is referred to as the conventional atmospheric lepton flux. In contrast, hadrons with much shorter lifetimes, such as charm hadrons, decay promptly over a wide energy range. The prompt atmospheric lepton flux from charm hadron decays scales as $\sim E_{\ell}^{-\gamma}$, up to lepton energies $E_{\ell}\sim 2\times 10^{7}$ GeV Gondolo et al. (1996). In addition to charm hadron decays, prompt muon production also receives contributions from electromagnetic decays of light unflavored mesons such as $\eta,\ \eta^{\prime},\ \rho^{0},\ \omega,$ and $\phi$, despite their small branching fractions ($\sim 10^{-5}$$10^{-4}$) Illana et al. (2011). Their extremely short lifetimes ensure decay before interaction, enhancing the prompt atmospheric muon flux at high energies by a factor of $\sim 2$ relative to the prompt neutrino flux.
@@
Figure 1: The atmospheric ($\mu^{+}+\mu^{-}$) flux and atmospheric ($\nu_{\mu}+\bar{\nu}_{\mu}$) flux, both scaled by $E^{3}$ and angle-averaged over zenith angles less than 60, from MCEq Fedynitch et al. (2019) using H3a cosmic ray flux model Gaisser (2012) and Sibyll-2.3c Riehn et al. (2018); Fedynitch et al. (2019) for interactions. The dashed lines show the prompt contributions, and the solid lines show the total contributions (conventional + prompt).
Prompt neutrinos from charm decays in the atmosphere are often considered an irreducible background to astrophysical neutrinos, since both fluxes are nearly isotropic at Earth. Their spectral behaviors differ, offering possible means of discrimination between the prompt atmospheric background and astrophysical neutrinos. In this paper, we focus on using the atmospheric muon flux as a complementary probe to atmospheric neutrinos. While many previous studies have focused primarily on neutrino fluxes Laha and Brodsky (2017); Halzen and Wille (2016); Giannini et al. (2018); Goncalves et al. (2022); Ostapchenko et al. (2023); Garzelli et al. (2024), our analysis expands to include both muons and neutrinos (see also, ref. Das et al. (2025)). As we describe in the next section, measurements of the atmospheric muon flux, particularly in the energy range where the prompt muon flux begins to dominate, exceed theory predictions, although with large experimental error bars Soldin et al. (2024); Soldin (2019); Aartsen and others (2016a). To solve this discrepancy, we first consider potential contributions from intrinsic charm. The intrinsic charm contributions are also constrained by prompt neutrino measurements. Therefore, with existing upper bounds on the prompt atmospheric muon neutrino flux from IceCube Aartsen and others (2016b, a); Abbasi and others (2025a), we also assess the extent to which the contribution of unflavored mesons to the prompt atmospheric muon flux may be enhanced.
The paper is organized as follows: in Section 2, we discuss the cascade equations that describe the development of the atmospheric lepton fluxes as cosmic rays and the hadrons they produce propagate through the atmosphere. We also show results from the cascade equations for muons and muon neutrinos, along with IceCube data and upper bounds on the prompt neutrino flux, respectively. In Section 3, we introduce the framework we use for incorporating intrinsic charm effects. Section 4 presents our results, followed by a discussion and our conclusions in Sections 5 and 6, respectively. For completeness, the Appendix includes details of the charm particle decays to leptons, needed to evaluate the lepton fluxes.
@@ 2.1 Cascade equations
Quantitative evaluations of the atmospheric flux begin with the incident all-nucleon cosmic ray spectrum at the top of the atmosphere, $\phi_{N}^{0}(E)$. The flux per solid angle $\phi_{N}^{0}$ is isotropic at high energies. The cascade equations that describe the development of particle fluxes with subsequent interactions and decays Lipari (1993); Gondolo et al. (1996). The cascade equations are evaluated as a function of column depth $X$ which depends on the atmospheric density ($\rho$), path length ($l$) traversed by the particle in the atmosphere, altitude ($h$) and zenith angle ($\theta$),
@@
The cascade equations can be solved approximately in the hadron-decay-dominated and hadron-interaction-dominated regimes Lipari (1993); Gondolo et al. (1996). This method, called the $Z$-moment method, approximates eq. (3) for production and decay processes by putting $\phi_{k}(E^{\prime},X)/\phi_{k}(E,X)\simeq\phi_{k}(E^{\prime},0)/\phi_{k}(E,0)$ so that
@@
An alternative approach for evaluating the atmospheric lepton fluxes is to solve the cascade equations directly. The Matrix Cascade Equations (MCEq) code Fedynitch et al. (2015, 2019) solves the cascade equations using a matrix approach. While the MCEq results are more precise, given hadronic production inputs, the $Z$-moment method is straightforward to implement. The two approaches differ by $5-20\%$ depending on energy and angle Gaisser et al. (2020). In this paper, we use the MCEq code with some modifications outlined below to evaluate all atmospheric lepton fluxes. Another approach is to use Monte Carlo air shower simulations. Recent work Neste et al. (2025) shows good agreement with MCEq.
@@ 2.2 Muon and muon neutrino fluxes from MCEq
MCEq is equipped with several hadronic interaction models and cosmic ray nucleon flux models. We use the Sibyll2.3c Riehn et al. (2018); Fedynitch et al. (2019) model for hadronic interactions, and the H3a cosmic ray flux and composition model Gaisser (2012). We discuss the impact of different hadronic models and all-nucleon cosmic ray flux inputs in Section 4.3.
@@
The implementation of charm production in Sibyll-2.3c follows a phenomenological model based on soft and hard minijets Fedynitch et al. (2019). For LHCb kinematic distributions for charmed-meson production, the overall combination of perturbative and non-perturbative contributions in Sibyll-2.3c is consistent with the next-to-leading order QCD predictions of refs. Gauld et al. (2015); Garzelli et al. (2017); Zenaiev and others (2015) that come from collinear factorization with partonic cross sections from perturbative QCD and non-perturbative contributions from both parton distribution functions and fragmentation functions/hadronization. While some components of the implementation of charm production in Sibyll-2.3c are described as “intrinsic” charm, the charm hadrons do not have the same energy distribution, shape and normalization, as in the model for intrinsic charm that we introduce in Section 3. Even though Sibyll-2.3c’s implementation of charm production has a weighted mix of perturbative and non-perturbative components, we will nevertheless refer to Sibyll-2.3c’s contribution to the prompt atmospheric lepton fluxes as coming from perturbative QCD (pQCD($c\bar{c}$)) to distinguish it from the intrinsic charm contribution we introduce in Section 3.
@@
Figure 2 shows the MCEq results for the separate contributions of pions, kaons and (prompt) charm hadrons to the muon (left) and muon neutrino (right) atmospheric fluxes for zenith angle $\theta_{\rm zen}=0^{\circ}$ (vertical). As noted above, the muon flux also includes contributions from light unflavored mesons to its prompt component. The standard release of MCEq does not include $\Lambda_{c}$ production and decay. We include perturbative $\Lambda_{c}$ production by extracting the $D^{0}$ production spectrum from MCEq and applying a constant suppression factor of 0.10 that reflects the relative charm quark fragmentation fractions to $\Lambda_{c}$ compared to $D^{0}$ in proton-proton collisions Lisovyi et al. (2016). The final results of our paper are not very sensitive to this suppression factor because charm contributions to the atmospheric $\mu^{+}+\mu^{-}$, $\nu_{\mu}+\bar{\nu}_{\mu}$ and $\nu_{e}+\bar{\nu}_{e}$ fluxes are dominated by the semi-leptonic decays of $D^{\pm},\ D^{0}$ and $\bar{D}^{0}$ mesons. For semi-leptonic $\Lambda_{c}$ decays, we follow refs. Bugaev et al. (1998) and Buras (1976) for the energy distributions of muons and neutrinos, which we have also described in the Appendix. While the details of the semi-leptonic $\Lambda_{c}$ decay distribution are not essential for the pQCD($c\bar{c})$ contributions to the prompt lepton fluxes, a $\Lambda_{c}$ is always produced in our implementation in intrinsic-charm induced events, as discussed in section 3.
@@
Figure 2 exhibits an high-energy behavior that reflects the dominance of the prompt contributions to the atmospheric lepton fluxes, visible in their harder spectra compared to the conventional fluxes. The plots show fluxes for vertical trajectories entering from the zenith, $\theta_{\rm zen}=0^{\circ}$. The zenith angle dependence is primarily in the conventional flux, where pions and kaons are more likely to interact than decay. The conventional flux increases with zenith angle according to $1/\cos\theta_{\rm zen}$ for $\theta_{\rm zen}\leq 70^{\circ}$ Klimushin et al. (2001). At very high energies, the prompt contributions will also vary with zenith angle, however, at the energies we consider for the rest of the paper, $E\lesssim 10^{6}$ GeV, the prompt atmospheric lepton fluxes are essentially isotropic.
@@
We can also obtain di-muon pairs through the process $\gamma\to\mu^{+}\mu^{-}$. There are two possible sources for this channel. The first is from diffuse cosmic rays interacting in the atmosphere and producing high energy photons, which can subsequently convert into muon pairs. However, this contribution is expected to be very small even at the highest $\gamma$-ray energies Illana et al. (2011). The second source is the diffuse $\gamma$-ray flux from the Galactic plane. According to ref. Cao and others (2023), at around 100 TeV the diffuse Galactic $\gamma$-ray flux is approximately $10^{-5}$ of the diffuse cosmic ray flux. Thus, diffuse Galactic $\gamma$-ray flux will also provide negligible contribution to the overall atmospheric muon flux.
Figure 3: Left: MCEq predictions for angle-averaged (up to zenith angle of $60^{\circ}$) conventional and prompt ($\mu^{+}+\mu^{-}$) flux, compared to the experimentally measured spectrum of high-energy muons using data for $\theta_{\rm zen}\leq 60^{\circ}$ from refs. Soldin et al. (2024); Soldin (2019); Aartsen and others (2016a). The unflavored and pQCD($c\bar{c}$) curves are nearly equal in the muon energy range shown in the figure. Right: Comparison of MCEq pQCD($c\bar{c}$) prompt ($\nu_{\mu}+\bar{\nu}_{\mu}$) flux with the IceCube upper limit on the prompt flux from Ref. Abbasi and others (2025a). The angle-averaged (up to zenith angle of $60^{\circ}$) conventional neutrino flux from MCEq is also shown. The prompt neutrino flux is isotropic in the energy range considered here. The cosmic ray all-nucleon flux model H3a and the hadronic interaction model Sibyll-2.3c are used as input of MCEq.